Hi! I'm really confused on a two-step equation. It says, "Nate's age is 3 years less than 2 times Ruth's age. Ruth is 8. How old is Nate?" The teacher says to write and equation and then solve it. I know that Nate's age is 13, because 8 * 2 - 3 equals 13. But I don't know how to write an equation that works? I've tried n / 2 + 3 = 8, n / 2 - 3 = 8, n / 3 + 2 = 8, n / 3 - 2 = 8, 2 * n + 3 = 8, 2 * n - 3 = 8, 3 * n + 2 = 8, 3 * n - 2 = 8.
Please help me if you can. xx

Answer by richwmiller(17219)

Nate's age is 3 years less than 2 times Ruth's age. Ruth is 8.
r=ruth
n=nate
n=2*r-3
r=8
n=2*8-3
16-3=n
n=13